OpenStudy (anonymous):

Find the properties for the ellipse with the equation 36x^2 + 4y^2 = 9. What does e=?

OpenStudy (anonymous):

sqrt(1-b^2/a^2)

OpenStudy (anonymous):

lets put the equation in standard form 36x^2 + 4y^2 = 9 divide by 9 throughout => 4x^2 + (4/9)y^2 = 1 => x^2/(1/2)^2 + y^2/(3/2)^2 = 1 a = 1/2 b = 3/2

OpenStudy (phi):

Defcon, the first thing you should do is put the equation in standard form divide both sides by 9 then make the coefficients of x^2 and y^2 the denominator

OpenStudy (anonymous):

Silence...

OpenStudy (anonymous):

c = sqrt(a^2-b^2) once you have c, you can find e using e = c/a

OpenStudy (anonymous):

Not necessary to calculate c

OpenStudy (anonymous):

or a, or b

OpenStudy (anonymous):

eccentricity is defined in terms of c its a ratio of distance from center and semi major axis length

OpenStudy (anonymous):

e = sqrt(1-b^2/a^2)

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

It is the same thing but you do not need to calculate c

OpenStudy (anonymous):

I got it now thanks guys